# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

### v1

@@ Abstract
The past few years have produced a series of spectacular advances in the decoding of speech from brain activity. The engine of these advances has been the acquisition of labelled data, with increasingly large datasets acquired from single subjects. However, participants exhibit anatomical and other individual differences, and datasets use varied scanners and task designs. As a result, prior work has struggled to leverage data from multiple subjects, multiple datasets, multiple tasks, and unlabelled datasets. In turn, the field has not benefited from the rapidly growing number of open neural data repositories to exploit large-scale data and deep learning. To address this, we develop an initial set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings. Experimental results show that representations learned with these objectives generalise across subjects, datasets, and tasks, and are also learned faster than using only labelled data. In addition, we set new benchmarks for two foundational speech decoding tasks. Taken together, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
Figure 1: Leveraging unlabelled data for speech decoding. We use abundant unlabelled neuroimaging data to pre-train a neural network using tasks that do not require explicit labels. The tasks apply a randomly selected transformation $T$ to the data and the network predicts the transformation. We fine-tune the pre-trained network on labelled data, achieving faster generalisation owing to the strength of the representation learned in pre-training.
@@ 1 Introduction
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet fully assimilated. State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive [40, 60] or non-invasive brain recordings [52], mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain. With few exceptions at present (e.g. [17]), speech decoding models tend not to train on data from more than one subject. Moreover, they do not combine data from multiple datasets and in general do not utilise unlabelled data, or data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations, as has begun to happen in neighbouring sub-fields (e.g. [27]), holds the promise of training models on collective, internet-scale data.
@@ 4.1 Experimental setup
Training. We pre-train all models to completion and then shallow or deep fine-tune on labelled data for each task, using this to analyse the generalisability of the pre-trained representation. Appendix B provides complete training details for all experiments.
@@ 4.2 Learning Generalisable Representations Using Pretext Tasks
To understand whether the pretext tasks produce useful representations, we investigate how they generalise independently with shallow fine-tuning (i.e. with the representation frozen). In Table 1, we show that all of our pretext tasks lead to better than chance accuracy on speech detection, while for voicing the band prediction task does best. In all cases, the tasks either match or beat a randomly initialised and shallow fine-tuned baseline. The results are statistically significant whenever we beat the baselines and provide initial evidence that all of our tasks are useful for learning representations that generalise downstream to speech decoding.
Pretext task Fine-tuned balanced accuracy
Speech $t$ $p$ Voicing $t$ $p$
None (random backbone) $0.500\pm 0.000$- -- $0.500\pm 0.000$- --
Phase($\rho=0.5$) $0.538\pm 0.002$ $18.5$ $0.001$ $0.500\pm 0.000$- --
Amplitude($\rho=0.2$) $\mathbf{0.555}\pm 0.002$ $23.3$ $0.001$ $0.500\pm 0.000$- --
Band $0.551\pm 0.005$ $10.0$ $0.005$ $\mathbf{0.503}\pm 0.001$ $4.1$ $0.03$
@@
Table 1: Pre-training with pretext tasks produces generalisable representations. We provide the test accuracy at the point of best validation accuracy (early stopping). In the none baseline, the network backbone is randomly initialised before fine-tuning (i.e. there is no pre-training). The uncertainty is the standard error of the mean over three random seeds. We quote the $t$-score and $p$-value for each result from a one-sample one-sided $t$-test where the population mean is chance-level accuracy. Models are trained on all subjects in Armeni et al. [2] and evaluated on held-out sessions.
@@
The strength of band prediction in voicing classification may be due to discriminating between theta waves and other frequency bands as theta waves are associated with phonetic feature processing. The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search (Appendix C). We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors. Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset [26]. In contrast, a larger $\rho$ for phase shift prediction may be best because phase is descriptive of neural synchrony which is distributed across the brain. Unlike localised responses, a large proportion of the sensors can detect this feature.
4.3 Scaling Speech Decoding With Unlabelled Data
Previously, we pre-trained and fine-tuned using data from Armeni et al. [2]. This setting was useful to understand whether our pretext tasks benefited the representation but did not demonstrate if they were effective in a real scarce data setting, where representations must generalise across datasets in order to scale. Furthermore, we did not investigate the scaling behaviour with downstream tasks. Here, we pre-train with Cam-CAN [48, 53], which does not have any speech decoding labels, and fine-tune for ten epochs on Armeni et al. [2] and Gwilliams et al. [24], showcasing transfer with minimal fine-tuning. This simulates the scenario in which models need to generalise quickly with very little calibration data from new patients in order to be deployed as a speech BCI. We use shallow fine-tuning for speech detection and deep fine-tuning for voicing. As evidenced by Table 1, we require the latter because voicing classification is a more difficult task, necessitating end-to-end training with more parameters to observe generalisation.
Figure 3: Scaling unlabelled data improves downstream generalisation. We pre-train the model on increasing amounts of unlabelled data and fine-tune on voicing and speech detection. No pre-training shows best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). We show only one no pre-training line as the result is the same for both datasets. The shaded area shows the standard error of the mean across three seeds.
We scaled up the pre-training dataset by increasing the number of subjects. Since this led to consistent and almost monotonic improvements in downstream accuracy, our self-supervised method is an exception to the common consensus that pooling subjects worsens generalisation. Moreover, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation. This is particularly notable as the Armeni et al. [2], Gwilliams et al. [24], and pre-training datasets all use different scanners entirely. Performing well across these datasets indicates that our pretext tasks successfully generate representations that are generalisable across heterogeneous scanners. Beyond hardware, our results also show improvements on distinct tasks, thus achieving cross-task generalisation. Finally, our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks, showing that remarkably any unlabelled data can be used to scale.
@@ 4.3 Scaling Speech Decoding With Unlabelled Data
Whilst we have shown that our representations generalise faster than without pre-training, do they outperform supervised baselines when trained to saturation? While voicing requires deep fine-tuning to see a substantial learning signal, the speech detection task, with which we use shallow fine-tuning, can be compared directly to similarly parameterised supervised classifiers. Table 2 provides this comparison. A representation pre-trained with 160 hours of unlabelled data significantly outperforms chance, a randomly initialised backbone, and supervised training on both downstream datasets. Here, the supervised classifier has significantly more parameters than the models that probe the network backbone because the input dimension is much larger without an encoder. Even with this bias towards the supervised classifier, the self-supervised case performs better. Thus, it is apparent that pre-training with unlabelled data significantly improves downstream generalisation.
@@
   Linear $0.5853\pm 0.0017$ $52$ $0.0003$ $0.5011\pm 0.0001$ $9$ $0.006$
   MLP (two layers) $\mathbf{0.6101}\pm 0.0017$ $51$ $0.0002$ $\mathbf{0.5230}\pm 0.0002$ $113$ $0.00004$
@@
Table 2: Self-supervised pre-training outperforms supervised classifiers. We use shallow fine-tuning on top of a randomly initialised and pre-trained backbone, comparing these to a supervised classifier trained to completion. We quote test accuracy at the point of best validation accuracy (early stopping) and show uncertainty as the standard error of the mean from three seeds. We calculate the $t$-score and $p$-value from one-sample one-sided $t$-tests where the population mean is chance.
4.4 Generic Self-Supervised Representations Improve Speech Decoding
This outcome indicates that our self-supervised losses are the main contributor to downstream accuracy. The MEG signal appears rich, encapsulating lots of information, with the speech component being a subset distributed across the representation space. Learning with specific labels produces representations that are too specialised to improve on other speech decoding tasks. This does not preclude there being some—or a combination of—labelled tasks that provide more general inductive biases for speech. However, the overall point remains salient. The generic property is lost with specific labels while it is well-preserved with SSL, providing another reminder of the bitter lesson.
@@ 4.4 Generic Self-Supervised Representations Improve Speech Decoding
4.5 Generalising Zero-Shot To Novel Subjects
@@
Brain data is variable across participants, leading to difficulty transferring models to novel subjects [11]. Whilst we have shown generalisation across subjects with our method, here, we investigate how well we can generalise to novel subjects—an even more difficult challenge. We pre-train then fine-tune our models on a subset of Gwilliams et al. [24], holding out three subjects for evaluation.
We conduct these experiments without conditioning, with subject embeddings, and with FiLM to see if they have any influence on subject generalisation. We represent embeddings as learnable vectors, unique to each subject, that are concatenated to the backbone representation. For FiLM, we use the same embeddings as inputs to the functions that modulate the backbone representation.
Subject Subject voicing balanced accuracy
conditioning Seen $t$ $p$ Unseen (zero-shot) $t$ $p$
None $0.5118\pm 0.001$ $8.4$ $0.007$ $0.5057\pm 0.001$ $4.5$ $0.02$
Embedding $0.5168\pm 0.0006$ $27.9$ $0.0006$ $0.5052\pm 0.0007$ $7.5$ $0.009$
Chance-level $0.5$- -- $0.5$- --
Table 3: Self-supervised representations generalise zero-shot to novel subjects. We quote test accuracy at the point of best validation accuracy (early stopping). Uncertainty is the standard error of the mean and we show the $t$-score and $p$-value from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
@@ Appendix A Supervised Heard Speech Decoding
Appendix B Experiment Details

### v2

@@ Abstract
The past few years have produced a series of spectacular advances in the decoding of speech from brain activity. The engine of these advances has been the acquisition of labelled data, with increasingly large datasets acquired from single subjects. However, participants exhibit anatomical and other individual differences, and datasets use varied scanners and task designs. As a result, prior work has struggled to leverage data from multiple subjects, multiple datasets, multiple tasks, and unlabelled datasets. In turn, the field has not benefited from the rapidly growing number of open neural data repositories to exploit large-scale data and deep learning. To address this, we develop an initial set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings. Experimental results show that representations learned with these objectives scale with data, generalise across subjects, datasets, and tasks, and are also learned faster than using only labelled data. In addition, we set new benchmarks for two foundational speech decoding tasks. Taken together, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
Figure 1: Leveraging unlabelled data for speech decoding. We pre-train a neural network using tasks that generate implicit labels from abundant unlabelled neuroimaging data, permitting learning from large heterogeneous datasets. The tasks apply a randomly selected transformation $T$ to the data and the network predicts the transformation. We fine-tune the pre-trained network on labelled data, achieving faster generalisation owing to the strength of the representation learned in pre-training.
@@ 1 Introduction
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated. State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive [40, 60] or non-invasive brain recordings [52], mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain. With few exceptions at present (e.g. [17]), speech decoding models tend not to train on data from more than one subject. Moreover, they do not combine data from multiple datasets and in general do not utilise unlabelled data, or data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations, as has begun to happen in neighbouring sub-fields (e.g. [27]), holds the promise of training models on collective, internet-scale data.
@@ 4.1 Experimental setup
Training. We pre-train all models to completion and then shallow or deep fine-tune on labelled data for each task, using this to analyse the generalisability of the pre-trained representation. Appendix E provides complete training details for all experiments.
@@ 4.2 Learning Generalisable Representations Using Pretext Tasks
We investigate whether our self-supervised objectives produce useful representations by first learning a representation with each pretext task independently and then analysing its downstream generalisation using shallow fine-tuning (i.e. with the representation frozen). In Table 1, we show that all of our pretext tasks lead to better than chance accuracy on speech detection, and critically, outperform a randomly initialised and shallow fine-tuned baseline. The results are statistically significant in all cases, providing initial evidence that all of our tasks are helpful in learning representations that generalise downstream to speech decoding.
Interestingly, the combination of all pretext tasks leads to better generalisation than any task on its own. As we hypothesised earlier, this may be because our pretext tasks capture complementary properties in time- and frequency-space, enforcing that our representation includes more salient features for speech decoding. For our other downstream task, voicing classification, we were not able to use shallow fine-tuning to generalise to a statistically significant degree with any of the pretext tasks nor with a random backbone. It may be too difficult to learn with shallow fine-tuning and so we turn to using deep fine-tuning for this task in the next section.
Pretext task Speech detection balanced accuracy
Armeni $t$ $p$ Gwilliams $t$ $p$
Amp($\rho=0.2$) $0.5846\pm 0.0032$ $26$ $7\mathrm{e}{-4}$ $0.5116\pm 0.0004$ $29$ $6\mathrm{e}{-4}$
Phase($\rho=0.5$) $0.5831\pm 0.0029$ $29$ $6\mathrm{e}{-4}$ $0.5093\pm 0.0014$ $7$ $1\mathrm{e}{-2}$
Band $0.5941\pm 0.0024$ $39$ $2\mathrm{e}{-3}$ $0.5177\pm 0.0028$ $6$ $1\mathrm{e}{-2}$
All $\mathbf{0.6011}\pm 0.0018$ $56$ $2\mathrm{e}{-4}$ $\mathbf{0.5206}\pm 0.0010$ $22$ $1\mathrm{e}{-3}$
None (rand. backbone) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
@@
Table 1: Pre-training with pretext tasks produces generalisable representations. We provide the test accuracy at the point of best validation accuracy (early stopping). In the none baseline, the network backbone is randomly initialised before fine-tuning (i.e. there is no pre-training). When all pretext tasks are used, their losses are weighted equally. The uncertainty is the standard error of the mean over three random seeds. We quote the $t$-score and $p$-value for each result from a one-sample one-sided $t$-test where the population mean is chance-level accuracy. Models are pre-trained on all of Cam-CAN [48, 53] and fine-tuned on Armeni et al. [2] or Gwilliams et al. [24] respectively.
@@
Among the pretext tasks, band prediction performs best. This could be because the representation easily identifies phase-locking to speech onset in theta waves [43]. However, further investigation is necessary to determine which frequency band is most informative. The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search (Appendix F). We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors. Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset [26]. In contrast, a larger $\rho$ for phase shift prediction may be best because phase is descriptive of neural synchrony which is distributed across the brain. Unlike localised responses, a large proportion of the sensors can detect this feature.
@@
Whilst we have shown that our representations generalise, do they outperform supervised baselines when trained to saturation? Using shallow fine-tuning once more, we compare our approach directly to similarly parameterised supervised classifiers in Table 2. A representation pre-trained with all pretext tasks significantly outperforms chance, a randomly initialised backbone, and supervised training on both downstream datasets. Here, the supervised classifier has significantly more parameters than the models that probe the network backbone because the input dimension is far larger without an encoder. Even with this bias favouring the supervised classifier, the self-supervised case performs better. Thus, it is apparent that pre-training with unlabelled data significantly improves downstream generalisation.
@@
   Linear $0.5853\pm 0.0017$ $52$ $3\mathrm{e}{-4}$ $0.5011\pm 0.0001$ $9$ $6\mathrm{e}{-3}$
   MLP (two layers) $\mathbf{0.6011}\pm 0.0018$ $56$ $2\mathrm{e}{-4}$ $\mathbf{0.5206}\pm 0.0010$ $22$ $1\mathrm{e}{-3}$
@@
Table 2: Self-supervised pre-training outperforms supervised classifiers. We use shallow fine-tuning on top of a randomly initialised and pre-trained backbone, comparing these to a supervised classifier trained to completion. We quote test accuracy at the point of best validation accuracy (early stopping) and show uncertainty as the standard error of the mean from three seeds. We calculate the $t$-score and $p$-value from one-sample one-sided $t$-tests where the population mean is chance. Self-supervised models are pre-trained on all of Cam-CAN [48, 53] and shallow fine-tuned on Armeni et al. [2] or Gwilliams et al. [24] while the supervised models are trained entirely on the latter datasets.
4.3 Scaling Speech Decoding With Unlabelled Data
Figure 3: Scaling unlabelled data improves downstream generalisation. We pre-train the model on increasing amounts of unlabelled data from Cam-CAN [48, 53] and fine-tune on voicing and speech detection on the Armeni et al. [2] and Gwilliams et al. [24] datasets. No pre-training shows best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). We show only one no pre-training line as the result is the same for both datasets. The shaded area shows the standard error of the mean across three seeds.
Previously, we showed that the combination of all pretext tasks leads to the best representations and that this outperforms a randomly initialised backbone as well as a supervised classifier. In this section, we examine how these pretext tasks fare as we increase the amount of unlabelled data, analysing scaling performance on downstream tasks. As before, we pre-train with the combined pretext tasks using Cam-CAN [48, 53], which does not have any speech decoding labels, and now fine-tune for ten epochs on Armeni et al. [2] and Gwilliams et al. [24], showcasing transfer with minimal fine-tuning. This simulates the scenario in which models need to generalise quickly with very little calibration data from new patients in order to be deployed as a speech BCI. Where earlier we analysed only speech detection, here, we also focus on voicing classification by using deep fine-tuning for this task only. As evidenced by Section 4.2, we require this because voicing classification is a more difficult task, necessitating end-to-end training with more parameters to observe generalisation.
We scaled up the pre-training dataset by increasing the number of subjects. Since this led to consistent and almost monotonic improvements in downstream accuracy, our self-supervised method is an exception to the common consensus that pooling subjects worsens generalisation. Moreover, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation. This is particularly notable as the Armeni et al. [2], Gwilliams et al. [24], and pre-training datasets all use different scanners entirely. Performing well across these datasets indicates that, together, our architecture and pretext tasks successfully generate representations that are generalisable across heterogeneous scanners.
Beyond hardware, our results also show improvements on distinct tasks, namely speech detection and voicing classification, thus illustrating cross-task generalisation. Thus, our pretext tasks are sufficiently generic to produce representations that generalise to multiple downstream tasks, indicating their applicability to speech decoding more generally. Finally, our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks. Remarkably, this shows that brain data collected from any task (including those that are not linguistic) can be used to scale downstream speech decoding performance.
@@ 4.3 Scaling Speech Decoding With Unlabelled Data
4.4 Scaling Unlabelled Data Improves Zero-Shot Generalisation To Novel Subjects
@@
Brain data is variable across participants, leading to difficulty transferring models to novel subjects [11]. Whilst we have shown generalisation across subjects with our method, here, we investigate how well we can generalise to novel subjects—an even more difficult challenge. This is critical in order to widely deploy speech BCIs for new patients. In the following experiment, we pre-train using the same setup as Figure 3, but fine-tune only on Gwilliams et al. [24]. We use FiLM conditioning as this performs best for seen and unseen subjects (Appendix B). During training, we hold out three subjects with which we evaluate novel subject generalisation (i.e. unseen subject performance).
Figure 4: Scaling pre-training data improves novel subject generalisation. As before, we pre-train with increasing amounts of unlabelled data from Cam-CAN [48, 53], but now fine-tune only on Gwilliams et al. [24]. In the seen case, we evaluate on held-out sessions from subjects that were used in the training data. For the unseen case we evaluate on three completely held-out subjects. No pre-training shows the best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). We show only one no pre-training line as the result is the same for both datasets. The shaded area is the standard error of the mean over three seeds.
Figure 4 shows that scaling up the amount of unlabelled data used in pre-training not only improves accuracy on subjects previously seen, but also demonstrates a positive trend in performance for unseen (novel) subjects. This is a promising result, indicating that scaling with our method is an encouraging direction for resolving the issues related to individual variance faced by prior work.
@@ Appendix A Supervised Heard Speech Decoding
Appendix B Subject Conditioning
To investigate subject conditioning approaches, we pre-train then fine-tune our models on a subset of Gwilliams et al. [24], holding out three subjects for evaluation. We conduct these experiments without conditioning, with subject embeddings, and with FiLM to see if they have any influence on subject generalisation. We represent embeddings as learnable vectors, unique to each subject, that are concatenated to the backbone representation. For FiLM, we use the same embeddings as inputs to the functions that modulate the backbone representation.
Subject Subject voicing balanced accuracy
conditioning Seen $t$ $p$ Unseen (zero-shot) $t$ $p$
None $0.5118\pm 0.001$ $8.4$ $0.007$ $0.5057\pm 0.001$ $4.5$ $0.02$
Embedding $0.5168\pm 0.0006$ $27.9$ $0.0006$ $0.5052\pm 0.0007$ $7.5$ $0.009$
Chance-level $0.5$- -- $0.5$- --
Table 4: Subject conditioning improves generalisation on seen subjects but not novel subjects. We quote test accuracy at the point of best validation accuracy (early stopping). Uncertainty is the standard error of the mean and we show the $t$-score and $p$-value from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
Appendix C Aggregating Unlabelled Datasets
To scale up unlabelled data further than individual studies, we must be able to combine many existing datasets. As a preliminary investigation, we combine two of the largest public MEG datasets: MOUS [47] and Cam-CAN [48, 53]. We investigate how pre-training with these combined datasets affects downstream performance using the same experimental setup as Figure 3.
Balanced accuracy
Labels Pre-training data (hours) Speech detection Voicing
Armeni MOUS (160 hours) $0.5736\pm 0.0016$ $0.5220\pm 0.0013$
Cam-CAN (159 hours) $\mathbf{0.5911}\pm 0.0009$ $\mathbf{0.5241}\pm 0.0004$
Cam-CAN + MOUS (319 hours) $0.5872\pm 0.0012$ $0.5236\pm 0.0016$
Gwilliams MOUS (160 hours) $0.5060\pm 0.0006$ $0.5175\pm 0.0008$
Cam-CAN (159 hours) $\mathbf{0.5118}\pm 0.0011$ $\mathbf{0.5196}\pm 0.0010$
Cam-CAN + MOUS (319 hours) $0.5116\pm 0.0017$ $0.5187\pm 0.0013$
Table 5: Combining two unlabelled datasets does not improve over the best study. We quote test accuracy at the point of best validation accuracy (early stopping). The uncertainty is the standard error of the mean calculated over three seeds.
The results in Table 5 suggest that combining these two datasets is not strictly better than the best dataset on its own. Pre-training on Cam-CAN alone leads to the best performance in all instances. This suggests that Cam-CAN is a better pre-training dataset than MOUS—a surprising result given that MOUS and Armeni et al. [2] share the same scanning hardware and MOUS includes data from subjects listening to speech while Cam-CAN has no such language tasks. This could be due to Cam-CAN being a cleaner dataset as we found artefacts can significantly affect learning. Combining Cam-CAN with MOUS performs marginally worse or the same as Cam-CAN alone. For all but one instance they are within the standard error of the mean of each other. The combined datasets always outperform MOUS on its own.
In general, further investigation is necessary to determine how datasets should be best aggregated. Increasing the number of datasets could enable the network to model the variance between them and eventually improve over the best singular dataset. This remains an open question as we were only able to study up to two datasets in this paper.
Appendix D Task Labels Are Hard To Scale; Generic SSL Scales Best
This outcome indicates that our self-supervised losses are the main contributor to downstream accuracy. The MEG signal appears rich, encapsulating lots of information, with the speech component being a subset distributed across the representation space. Learning with specific labels produces representations that are too specialised to improve on other speech decoding tasks. This does not preclude there being some—or a combination of—labelled tasks that provide more general inductive biases for speech; nor does it mean more labelled data will not scale. However, the point remains salient. Labelled data is more difficult to scale given its scarcity and the generic property is lost with specific labels while it is well-preserved with SSL, providing another reminder of the bitter lesson.
Appendix E Experiment Details